OpenStudy (anonymous):

lim [2x^2/3x+1]^(-3x^2+2/5x-3)=0 x->∞

ganeshie8 (ganeshie8):

hey still here ?

OpenStudy (anonymous):

Yes Im here

ganeshie8 (ganeshie8):

\[\lim\limits_{x\to\infty}~\left(\dfrac{2x^2}{3x+1}\right)^{\left(\dfrac{-3x^2+2}{5x-3}\right)}\] is this the expression ?

OpenStudy (anonymous):

Yes, it is

OpenStudy (dan815):